Allison:

What percent is 1.6/2?

Pixel:

something over 50%

Allison:

I know.

Pixel:

maybe 56%

Hero:

You can take the percent of a number. Based on questions you've submitted in the past, I would rather you submit a screenshot of the way it is presented from your instructor before I offer any help with this one.

Allison:

Oh, it's a grade I got on a test, and I do not know how much that is in percentage.

Pixel:

oh

Pixel:

1.6/2 = 86%-96%

Allison:

Hmm.

dude:

Pixel thats close Divide 1.6 by 2 first

Pixel:

E.E

Allison:

0.8

dude:

Then multiply by 100

Pixel:

Dood i know my percent grades,

Allison:

80%?

dude:

Yes

Pixel:

e.e

Allison:

Huh, not bad o:

Allison:

Okee and 3.3/4? e.e

dude:

Divide 3.3 by 4

Pixel:

83% or 82.5

dude:

Right

Allison:

82.5%

Allison:

Derick lemme do it damnit

Pixel:

fine e.e

Pixel:

I don't help anybody anyway

Allison:

Rip e.e

Pixel:

because no one believes me

Allison:

No one said you're not believed.

Hero:

\(\dfrac{1.6}{2} = \dfrac{1 + \dfrac{6}{10}}{2} = \dfrac{\dfrac{10}{10} + \dfrac{6}{10}}{2} = \dfrac{\dfrac{16}{10}}{2} = \dfrac{16}{10}\times \dfrac{1}{2} = \dfrac{16}{20} = \dfrac{4}{5} = \dfrac{80}{100} = 80\%\)

Allison:

Holy Pluto.

Allison:

Okee okee, I see you